### v2

@@ 3.2 Algebra of observables
We consider the algebra of observables generated by the spin-$s$ Pauli operators $X_{i}^{(s)}$, $Y_{i}^{(s)}$, and $Z_{i}^{(s)}$, defined in Section 6.1, along with the identity operator on the corresponding Hilbert space [^6] Given this initial set of operators, one can construct the associated von Neumann algebra by including limits of sequences of operators that converge in the weak operator topology..
Since we will be interested in the spectral properties of operators and their connection with free probability predictions, we will focus on a map of the form $\varphi(\cdot)=\mathbb{E}\,\text{Tr}(\cdot)/N$, which corresponds to an ensemble average of a infinite temperature thermal state.
4 DOS from free probability
@@ 4.1 Cauchy and R-transforms
In this section, we outline a procedure for approximating the DOS of the RP model (20) using tools from free probability theory. In the ergodic phase, corresponding to small values of $\gamma$, the DOS is well approximated by the Wigner semicircle distribution. As $\gamma$ increases and the system enters the fractal phase, the DOS gradually deforms, interpolating between the semicircular shape and a Gaussian profile. In the localized phase, which emerges at large $\gamma$, the DOS is well described by a Gaussian distribution. See Fig. 3 for a schematic illustration of this behavior across the three phases. Here, we are interested in analytically computing the DOS specifically for the fractal phase of the RP model. [^7] For other approaches to the calculation of the DOS of the RP model, we refer to kreynin1995density ; Venturelli:2022hka .
From the functional inverse of the resolvent of an operator $\mathcal{O}$ (i.e., the Cauchy transformation of the density of eigenvalues of the operator $\mathcal{O}$), we define the $R$-transformation
$G_{\mathcal{O}}^{-1}(z)\equiv\mathcal{B}_{\mathcal{O}}(z)=\frac{1}{z}+R_{\mathcal{O}}(z)~{}.$ (24)
We note that the expression for the resolvent of the operator $A$ in terms of its DOS, $\rho_{\mathcal{O}}(\lambda)$ as
@@
Here, the integral is over the support of the distribution $\rho_{\mathcal{O}}(\lambda)$. Note that $G_{\mathcal{O}}(z)$ is essentially the Cauchy transformation of $\rho_{\mathcal{O}}(\lambda)$.
When the operator $\mathcal{O}$ is the Hamiltonian of some quantum mechanical system belonging to some random matrix ensemble, one can relate the resolvent with the time evolution operator as follows. Consider the trace of the ensemble-averaged time evolution operator,
$\mathcal{U}_{H}(t)=\frac{1}{N}\mathbb{E}\big{(}\text{Tr}~{}e^{-iHt}\big{)}~{},$ (26)
which is related to the DOS through a Fourier transformation
$\rho_{H}(\lambda)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\text{d}t~{}e^{-it\lambda}~{}\mathcal{U}_{H}(t)~{}.$ (27)
Now, from the definition of the resolvent in (25), it is easy to see that $G_{H}(z)$ is related to the averaged time evolution operator as BREZIN1996697
$G_{H}(z)=i\int_{0}^{\infty}~{}\text{d}t~{}\mathcal{U}_{H}(t)~{}e^{-itz}~{}.$ (28)
@@ 4.2 Addition of two free operators: subordination formulae for the resolvent
Now, for addition of two such operators $P$ and $Q$, which are free, from the additivity of the $R$ transformation, we have, for the operator $C=P+Q$, the following,
$\mathcal{B}_{C}(z)=R_{C}(z)+\frac{1}{z}\,~{}~{}\rightarrow~{}~{}\mathcal{B}_{A}(z)=\mathcal{B}_{C}(z)-R_{B}(z)\,.$ (29)
Substituting $z\rightarrow G_{C}(z)$ in the argument of the last expression above, we have
$\mathcal{B}_{P}\big{(}G_{C}(z)\big{)}=z-R_{Q}\big{(}G_{C}(z)\big{)}\,.$ (30)
Finally, applying $G_{P}$ on both sides, we get a self-consistency equation for $G_{C}(z)$,
@@
One can perform an entirely analogous procedure for the operator $Q$ and obtain
@@
In the following, we use these formulae to obtain an expression for the DOS for the RP model. The choice between the two formulae for obtaining the density of eigenvalues of $C$ – which can be extracted from $G_{C}(z)$ using the Stieltjes inversion formula – depends on whether the resolvent and the $R$-transformation are more straightforward for the operators $P$ or $Q$.
Before proceeding further, we note that a self-consistency equation analogous to the one in (31) for the resolvent of the RP model was obtained in Venturelli:2022hka using the replica trick. The authors also discussed the connection of their formula with the one obtained from the law of addition in free probability. The discussion on the approximate evaluation of the DOS of the RP model below follows closely the discussion of Ref. Venturelli:2022hka . However, we have kept the discussion below general enough so that it applies not only for the RP model, but addition of two non-commutative operators, one of which has the Wigner semicircle as the DOS.
@@ 4.3 A perturbative scheme for the resolvent
For most of the choices of the operators $P$ and $Q$, the expressions for the corresponding resolvents are sufficiently complicated so that it is difficult to solve either of the equations in eqs. (31) or (32) for $G_{C}(z)$. Therefore, we now proceed as follows to obtain a perturbative expression for the DOS of an operator of the form $C=P+\alpha Q$, where $P$ and $Q$ are mutually free and the expressions for their DOS are assumed to be known. Here $\alpha=N^{-\gamma/2}$ is a parameter which we assume to be small so that one can obtain a perturbation series. Using the scaling transformation rule of the $R$-transformation, [^8] The $R$-transformation of the density of eigenvalues of an operator $P$ follows the scaling law: $R_{\alpha P}(z)=\alpha R_{P}(\alpha z)$, where $\alpha$ is a constant. and the definition (in Eq. (25)), we have, from Eq. (31),
@@
Here we have defined the characteristics function $\chi_{P}(u)$ of the distribution $\rho_{P}(\lambda)$ by a Fourier transformation
$\chi_{P}(u)=\int\text{d}\lambda~{}\rho_{P}(\lambda)~{}e^{-i\lambda u}~{}.$ (34)
@@
Next, expanding the exponential factor containing $R_{B}$ in Eq. (4.3) we have the following series expansion for $G_{C}(z)$,
@@
Writing in terms of the free cumulants ($\kappa_{j}(Q)$) of the operator $Q$, this has the form
$G_{C}(z)=\sum_{n}\frac{i^{n+1}\alpha^{n}}{n!}\Big{(}\sum_{j=1}^{\infty}\kappa_{j}(Q)\big{(}\alpha G_{C}(z)\big{)}^{j}\Big{)}^{n}\int_{0}^{\infty}~{}\text{d}u~{}\chi_{P}(-u)~{}u^{n}~{}e^{-iuz}\,.$ (36)
So far, this expression for $G_{c}(z)$ is exact. For a general operator $Q$, its $R$-transformation can be an arbitrary function of $z$, so that it is difficult to obtain a perturbation series with $\alpha$ as a small parameter. However, one particular case where the above expression becomes easily tractable is when the operator $B$ has Wigner semicircle as the DOS (which is the case also for the RP model considered in this paper), i.e., $\rho_{Q}(\lambda)=\frac{1}{2\pi}\sqrt{4-\lambda^{2}}$. In that case [^9] The Wigner semicircle is the analogue of the Gaussian distribution for non-commutative probability, i.e., this distribution has only the first free cumulant non-zero. Since $R(z)$ is the generator of the free cumulants, this is equal to $R(z)=z$ for the semicircle distribution. $R(z)=z$, and the expression in Eq. (35) (or (36)) simplifies to
@@
where to obtain the last expression, we have used an analogous relation to that of the one in (28), valid for the operator $P$.
@@
Therefore, all orders of correction can be obtained in terms $G_{C}^{(0)}(z)=G_{P}(z)$ (and its various order derivatives with respect to $z$), which in tern can be obtained from the characteristics function of the distribution $\rho_{P}(\lambda)$. Note also that due to the specific scaling law for the $R$-transformation, the $n$-order correction to $G_{C}^{(0)}(z)$ is proportional to $\alpha^{2n}$. The approximate DOS of the operator $C$ can now be obtained from $G_{C}(z)$ using the Stieltjes inversion formula (7),
@@
Figure 4: DOS of the RP model: comparison of approximate analytical expression with numerical results. The blue and black-dashed curves respectively represent the analytical DOS with first and second-order corrections. The red curve represents the Gaussian eigenvalue distribution of the matrix $A$. Here we have set $s=1/2,~{}\gamma=1.1,L=10$, and the histogram shows the numerical DOS averaged over 5000 independent realisations of the RP model Hamiltonian.
@@ 4.4 Approximate DOS of the RP model
Until now, the approximate expressions for $G_{C}(z)$ we have derived above are valid for any operator $P$ with a well-defined DOS, which is then added with an operator $Q$ having a semicircle density. Now specializing to the case of the RP model considered in this paper, we see that the eigenvalue distribution for the diagonal matrix $P\equiv A$ is a Gaussian distribution with zero mean and variance $\sigma^{2}=4/N$, and $Q\equiv B$. Hence, denoting the operator $C$ as the Hamiltonian $H$ of the RP model, we have
@@
[…]
In Figure 4, we have shown the DOS of the RP model obtained using this procedure along with the one obtained numerically. [^11] In Ref. Truong:2016wcq , an implicit formula for the DOS of a Hamiltonian where a random matrix perturbation ($V$) is added to a random (or a non-random matrix) matrix $H_{0}$ was derived using supersymmetry techniques, including the case of the RP model. The numerical results reported there seem to be consistent with the ones reported here. As can be seen, apart from a small region at the center of the spectrum, the approximate expression has a good agreement with numerical results.
@@ 5.1 Choice of operators
The $n$-point correlation functions of an operator $\mathcal{O}$ with appropriate normalization that we are interested in are defined as
@@
where the spin system consists of $L$ lattice sites, with $Z_{k}^{(s)}$ and $I_{k}$ being the Pauli $z$ matrices (of spin $s$) and identity operators acting on site $k$. The parameter $\epsilon$ distinguishes between spin types: $\epsilon=2$ for half-integer spins and $\epsilon=1$ for integer spins.
@@
Using the above definition, we study the behavior of the $n$-point functions
$F_{2n,(s)}^{(\gamma)}(t)=\frac{1}{(2s+1)^{L}}\mathbb{E}\,\mathrm{Tr}\big{[}(S_{1}(0)S_{1}(t))^{n}\big{]}=\langle(S_{1}(0)S_{1}(t))^{n}\rangle\,,~{}~{}n=0,1,2,\cdots\,,$ (50)
throughout the regime $\gamma\in[0,1]$, i.e., in the ergodic phase. Similarly, the evolution can be defined for $S_{L}$. From now on, we concentrate on operators with spin $s=1/2$ only, with different system sizes $L$, and omit the subscript $s$.
5.2 Approximate analytical expressions for the two-point function
In random matrix theory, one can show that the two-point functions of operators are closely related to the spectral form factor (SFF), while the higher-order correlation functions are also related to the higher-order SFFs  Cotler:2017jue . At early times, the SFF is well-approximated by the square of the partition function – corresponding to the disconnected contribution of the correlation function. We expect this relation to hold approximately in the ergodic phase of the RP model, particularly for small values of $\gamma$ [^12] Note that the relation between the two-point function and the SFF that was derived in Cotler:2017jue is only valid for random matrix ensembles having unitary invariance. However, for the RP mode, the rotational invariance is broken, even for small values of $\gamma$. Therefore, the relation in (51) between the two-point function and the disconnected part of the SFF that we propose below is essentially an observation. Following the approach of Cotler:2017jue , we estimate the early-time behavior of the two-point function as:
$F_{2}^{(\gamma)}(t)=\langle S_{1}(0)S_{1}(t)\rangle\approx Z(t)^{2}/N^{2}\,,\qquad Z(t)=N\int\textrm{d}E\,\rho(E)e^{-iEt}\,,\qquad(t\lesssim t_{d})$ (51)
where $t_{d}$ denotes the timescale up to which the approximation remains valid, $N$ is the rank of the Hamiltonian, and $\rho(E)$ is the DOS.
Figure 5: Behavior of 2-point (left), 4-point (middle), and $2n$-point functions of the operator $S_{1}$ (with $s=1/2$) in the RP model. The black dashed line indicates the analytic result (53) for $\gamma\leq 0.5$, which reduces to (54) for $\gamma=0$. We choose $L=8$, and a single realization of the Hamiltonian is taken.
Deep in the ergodic phase of the RP model (with sufficiently small value of the parameter $\gamma$), and sufficiently large $N$, the normalized DOS is well approximated by a Wigner semicircle kravtsov2015random ; guhr1996transitions

### v3

@@ 3.2 Algebra of observables
We consider the algebra of observables generated by the spin-$s$ Pauli operators $X_{i}^{(s)}$, $Y_{i}^{(s)}$, and $Z_{i}^{(s)}$, defined in Section 6.1, along with the identity operator on the corresponding Hilbert space. [^8] Given this initial set of operators, one can construct the associated von Neumann algebra by including limits of sequences of operators that converge in the weak operator topology. Since we will be interested in the spectral properties of operators and their connection with free probability predictions, we will focus on a map of the form $\varphi(\cdot)=\mathbb{E}\,\text{Tr}(\cdot)/N$, which corresponds to an ensemble average of a infinite temperature thermal state.
4 Density of states (DOS) from Free Probability
In this section we outline a procedure for approximating the DOS of the RP model (20) using tools from free probability theory. In the ergodic phase, corresponding to small values of $\gamma$, the DOS is well approximated by the Wigner semicircle distribution. As $\gamma$ increases and the system enters the fractal phase, the DOS gradually deforms, interpolating between the semicircular shape and a Gaussian profile. In the localized phase, which emerges at large $\gamma$, the DOS is well described by a Gaussian distribution. See Fig. 3 for a schematic illustration of this behavior across the three phases. Here, we are interested in analytically computing the DOS specifically for the fractal phase of the RP model. [^9] For other approaches to the calculation of the DOS of the RP model, we refer to kreynin1995density ; Venturelli:2022hka .
@@ 4.1 Cauchy and R-transforms
We recall that the functional inverse of the trace of the resolvent (i.e., the Cauchy transformation of the density of eigenvalues), and the $R$-transformation of an operator $\mathcal{O}$ is given by speicher2019lecturenotesfreeprobability ; nica2006lectures
$G_{\mathcal{O}}^{-1}(z)\equiv\mathcal{B}_{\mathcal{O}}(z)=\frac{1}{z}+R_{\mathcal{O}}(z)\,$, (24)
with the expression for the resolvent in terms of its DOS $\rho_{\mathcal{O}}(\lambda)$ as
@@
When the operator $\mathcal{O}$ is the Hamiltonian of a quantum mechanical system belonging to a random matrix ensemble, one can relate the resolvent to the time evolution operator as follows. Consider the trace of the ensemble-averaged time evolution operator:
$\mathcal{U}_{H}(t)=\frac{1}{N}\mathbb{E}\big(\text{Tr}~e^{-iHt}\big)\,$, (26)
which is related to the DOS through a Fourier transformation BREZIN1996697
$\rho_{H}(\lambda)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\text{d}t~e^{-it\lambda}~\mathcal{U}_{H}(t)\,.$ (27)
From the definition of the resolvent in (25), it is easy to see that $G_{H}(z)$ is related to the averaged time evolution operator as BREZIN1996697
$G_{H}(z)=i\int_{0}^{\infty}~\text{d}t~\mathcal{U}_{H}(t)~e^{-itz}\,.$ (28)
@@ 4.2 Addition of two free operators: subordination formulae for the resolvent
Let us consider two free operators $P$ and $Q$ which have well-defined DOS in the limit $N\rightarrow\infty$, and we are interested in the DOS of the sum of the operators $C=P+Q$. From the additivity of the $R$ transformation: $R_{C}=R_{P}+R_{Q}$ speicher2019lecturenotesfreeprobability ; nica2006lectures ; voiculescu1986addition , we write
$\mathcal{B}_{C}(z)=R_{C}(z)+\frac{1}{z}\,,~~\Rightarrow~~\mathcal{B}_{P}(z)=\mathcal{B}_{C}(z)-R_{Q}(z)\,.$ (29)
Substituting $z\rightarrow G_{C}(z)$ in the argument of the last expression above, we find
$\mathcal{B}_{P}\big(G_{C}(z)\big)=z-R_{Q}\big(G_{C}(z)\big)\,$, (30)
where we have used $\mathcal{B}_{C}(G_{C}(z))=z$, since $G_{C}$ and $\mathcal{B}_{C}$ are functional inverses to each other. Further applying $G_{P}$ on both sides and using the same functional inverse relationship, we get a self-consistency equation for $G_{C}(z)$:
@@
One can perform an entirely analogous procedure for the operator $Q$ to obtain
@@
In the following, we use these formulae to obtain an expression for the DOS for the RP model. The selection between the above two formulae (31) and (32) for determining the eigenvalue density of CC–which can be extracted from $G_{C}(z)$ through the Stieltjes inversion formula–is guided by the relative simplicity of applying the resolvent and the $R$-transformation to either operator $P$ or $Q$.
Before proceeding further, we observe that a self-consistency equation, analogous to the one in (31) for the resolvent of the RP model, was derived in Venturelli:2022hka through the replica trick. The authors also discussed the relationship between their formula and that obtained via the law of addition in free probability. The following discussion, which addresses the approximate evaluation of the DOS for the RP model, closely parallels the treatment in Ref. Venturelli:2022hka . Nevertheless, the exposition has been kept sufficiently general to encompass not only the RP model but also the addition of two non-commutative operators, one of which exhibits a Wigner semicircle as its DOS.
@@ 4.3 A perturbative scheme for the resolvent
For most choices of the operators $P$ and $Q$, the expressions for their respective resolvents are sufficiently intricate, making it challenging to solve either of the equations in Eqs. (31) or (32) for $G_{C}(z)$. Consequently, we proceed to derive a perturbative expression for the DOS of an operator of the form $C=P+\alpha Q$, where $P$ and $Q$ are mutually free, and their respective DOS expressions are assumed to be known. Here, $\alpha$ is a parameter which we assumed to be small, thereby allowing us to develop a perturbation series for the Cauchy transformation of the DOS of the operator $C$. For the RP model Hamiltonian, we have $\alpha=N^{-\gamma/2}$.
Utilizing the scaling transformation rule of the $R$-transformation, [^10] The $R$-transformation of the density of eigenvalues for an operator $P$ obeys the scaling law: $R_{\alpha P}(z)=\alpha R_{P}(\alpha z)$, where $\alpha$ is a constant. along with the definition given in Eq. (25), we obtain, from Eq. (31):
@@
Here we have defined the characteristic function $\chi_{P}(u)$ of the distribution $\rho_{P}(\lambda)$ by a Fourier transformation
$\chi_{P}(u):=\int\text{d}\lambda~\rho_{P}(\lambda)~e^{-i\lambda u}\,.$ (34)
@@
Next, expanding the exponential factor containing $R_{Q}$ in Eq. (4.3) we have the following series expansion for $G_{C}(z)$:
@@
Now, we express the $R$-transform of any operator in terms of free cumulants $\kappa_{j}$, defined by the following relation mingo :
$R(z):=\sum_{j=1}^{\infty}\kappa_{j}z^{j-1}\,.$ (36)
Hence, expressing in terms of the free cumulants $\kappa_{j}(Q)$ of the operator $Q$, $G_{C}(z)$ has the following generic form
$G_{C}(z)=\sum_{n}\frac{i^{n+1}\alpha^{n}}{n!}\Big(\sum_{j=1}^{\infty}\kappa_{j}(Q)\big(\alpha G_{C}(z)\big)^{j-1}\Big)^{n}\int_{0}^{\infty}~\text{d}u~\chi_{P}(-u)~u^{n}~e^{-iuz}\,.$ (37)
So far, the above expression for $G_{c}(z)$ is exact. For any generic operator $Q$, its $R$-transformation can be an arbitrary function of $z$, so that it is difficult to obtain a perturbation series with $\alpha$ as a small parameter. However, one particular case where the above expression becomes easily tractable is when the operator $Q$ has Wigner semicircle as the DOS (which is the case also for the RP model considered in this paper), i.e., $\rho_{Q}(\lambda)=\frac{1}{2\pi}\sqrt{4-\lambda^{2}}$. In that case [^11] The Wigner semicircle is the analogue of the Gaussian distribution for non-commutative probability. Its $R$-transform takes a particularly simple form: $R(z)=z$. Given that $R(z)$ generates free cumulants via the relation (36), the semicircle distribution is characterized by a single non-vanishing free cumulant, which is the second free cumulant and for the semi-circle distribution above, it has a value of unity. $R(z)=z$, and the expression in Eq. (35) (or (37)) simplifies to
@@
where to obtain the last expression, we have used an analogous relation to that of the one in (28), valid for the operator $P$, i.e., $G_{P}(z)=i\int_{0}^{\infty}\chi_{P}(-u)\,e^{-iuz}$, with $\chi(u)$ given by Eq. (34).
@@
Hence, all correction terms can be expressed in terms of $G_{C}^{(0)}(z)=G_{P}(z)$ and its derivatives with respect to $z$. These, in turn, can be derived from the characteristic function of the distribution $\rho_{P}(\lambda)$. Furthermore, owing to the specific scaling behavior of the $R$-transformation, the $n$-th order correction to $G_{C}^{(0)}(z)$ is proportional to $\alpha^{2n}$. The approximate DOS for the operator $C$ can then be determined from $G_{C}(z)$ using the Stieltjes inversion formula (7)
@@
Figure 4: DOS of the RP model: comparison of approximate analytical expression with numerical results. The blue and black-dashed curves respectively represent the analytical DOS with first and second-order corrections. The red curve represents the Gaussian eigenvalue distribution of the matrix AA. Here we have set $s=1/2,~\gamma=1.1,L=10$, and the histogram shows the numerical DOS averaged over 5000 independent realisations of the RP model Hamiltonian.
@@ 4.4 Approximate DOS of the RP model
Thus far, the derived approximate expressions for $G_{C}(z)$ hold for any operator $P$ with a well-defined DOS, combined with an operator $Q$ characterized by a semicircular density. Turning now to the specific case of the RP model examined in this paper, the eigenvalue distribution of the diagonal matrix $P\equiv A$ follows a Gaussian distribution with zero mean and variance $\sigma^{2}=4/N$, while $Q\equiv B$. Consequently, we identify the operator $C$ as the Hamiltonian $H$ of the RP model. Using the Gaussian DOS of $P$ as $\rho_{P}(\lambda)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{\lambda^{2}}{2\sigma^{2}}}$, we find $\chi_{P}(u)=e^{-\frac{1}{2}u^{2}\sigma^{2}}$, leading to the following zeroth-order resolvent of operator $C$:
@@
where $D(x)$ denotes the Dawson function: $D(x)=e^{-x^{2}}\int_{0}^{x}e^{t^{2}}~\text{d}t$. Using this expression for $G_{C}^{(0)}(z)$ in Eqs. (40) and (4.3), we obtain the first and second-order corrections to $G_{C}(z)$, which, in this scenario, can also be expressed in terms of the Dawson function. Consequently, the DOS of the RP model can be obtained from (42). It is important to note that, for the RP model, this approximation for the DOS holds when $\gamma>1$. [^12] Even though our formalism is general, the formula above captures only small corrections to the Gaussian distribution, which is characteristic of the localized phase. Naturally, such corrections might still provide an accurate description of the DOS in the fractal phase–an observation consistent with our findings. For $\gamma\lesssim 1$, however, the DOS resembles a deformed semicircle, and we do not expect the formula to yield quantitatively accurate predictions in this regime.
In Figure 4, we have shown the DOS of the RP model obtained using this procedure along with the one obtained numerically. [^13] In Ref. Truong:2016wcq , an implicit formula for the DOS of a Hamiltonian where a random matrix perturbation ($V$) is added to a random (or a non-random matrix) matrix $H_{0}$ was derived using supersymmetric techniques, including the case of the RP model. The numerical results reported there seem to be consistent with the ones reported here. As can be seen, the perturbative expression, even at $\mathcal{O}(\alpha^{2})$, has an excellent agreement with numerical results.
@@ 5.1 Choice of operators
The $2n$-point correlation functions of an operator $\mathcal{O}$ with appropriate normalization that we are interested in are defined as
@@
where the spin system consists of $L$ lattice sites, with $Z_{k}^{(s)}$ and $I_{k}$ being Pauli $\sigma_{z}$ matrix (of spin $s$) and identity operators acting on site $k$. The parameter $\epsilon$ distinguishes between spin types: $\epsilon=2$ for half-integer spins and $\epsilon=1$ for integer spins.
@@
Using the above definition, we study the behavior of the $2n$-point OTOCs:
$F_{2n,(s)}^{(\gamma)}(t)=\frac{1}{(2s+1)^{L}}\mathbb{E}\,\mathrm{Tr}\big[(S_{i}(0)S_{j}(t))^{n}\big]=\langle(S_{i}(0)S_{j}(t))^{n}\rangle\,,~~n=0,1,2,\cdots\,$, (51)
throughout the regime $\gamma\in[0,1]$, i.e., in the ergodic phase. From now on, we concentrate on operators with spin $s=1/2$ and omit the subscript $s$.
5.2 Correlation functions in the ergodic phase from asymptotic freeness
Using the fact that, deep in the ergodic phase of the RP model, the operator dynamics resembles that of Haar-random evolution, we can exploit the notion of freeness between random and deterministic matrices in the large-dimension limit to derive explicit expressions for the $2n$-point correlation functions. First, consider the two-point correlation function $\langle S_{i}(0)S_{i}(t)\rangle$, which can be expanded as
$\langle S_{i}(0)S_{i}(t)\rangle=\sum_{k,l}\frac{(it)^{k}(-it)^{l}}{k!~l!}\langle{S_{i}H^{k}S_{i}H^{l}}\rangle\,.$ (52)
To evaluate the mixed moment $\langle S_{i}H^{k}S_{i}H^{l}\rangle$, we use the fact that, deep in the ergodic phase, the Hamiltonian $H$ can be treated as free from the deterministic operator $S_{i}$ in the limit $N=(2s+1)^{L}\rightarrow\infty$. Since $S_{i}$ is traceless, the only non-vanishing contribution in the moment expansion comes from the diagram that factorizes as $\langle{S_{i}H^{k}S_{i}H^{l}}\rangle~=\langle{S_{i}S_{i}}\rangle\langle H^{k}\rangle\langle H^{l}\rangle$ Camargo:2025zxr .
Next, to calculate moments of the Hamiltonian of different orders, we need the DOS of the Hamiltonian. To this end, note that, deep in the ergodic phase of the RP model (with sufficiently small values of the parameter $\gamma$) and for large $N$, the normalized DOS is well approximated by a Wigner semicircle kravtsov2015random ; guhr1996transitions :